Determine the charge of each atom if it is composed of the following particles:


1. 5 electrons and 5 protons



2. 6 protons and 5 electrons



1
Expert's answer
2022-02-11T08:47:17-0500

The net charge of atom containing:

1. 5 electrons and 5 protons

"Q=5(-e)+5e=0"

2. 6 protons and 5 electrons

"Q=5(-e)+6e=e=1.6*10^{-19}\: \rm C"